OpenStudy (anonymous):

Can you help me solve? 10/3x + 4/3 = 7 + x/2x

OpenStudy (anonymous):

First find common denominators to make each side a single fraction, and then cross multiply. Alternatively, multiply each side by \(x\).

OpenStudy (cwrw238):

lcm of 3x,3 and 2x is 6x multiply through by 6x: 20 + 8x = 42x + 3x 37x = 20 x = 20/37

OpenStudy (anonymous):

is it x= 17/5x?

OpenStudy (anonymous):

cwrw238, But what does 20/37 have in common? Because the multiple choices are: x= 1/3 x= 17/5x x= 1/5 x= 1/6 (is this the answer)?

OpenStudy (cwrw238):

lets make sure i've have got the correct question maybe the right side of equation is (7+ x)/2x?

OpenStudy (cwrw238):

the way its been written is ambiguous

OpenStudy (anonymous):

yes, 10/3x + 4/3 equals 7 + x/ 2x

OpenStudy (cwrw238):

ok so 20 + 8x = 21 + 3x 5x = 1 x = 1/5

OpenStudy (anonymous):

thank you so much ! You were very helpful while still having the patience

OpenStudy (cwrw238):

its best to write it as (x+7) to avoid confusion

OpenStudy (cwrw238):

lol - i have my grandfathers patience

OpenStudy (anonymous):

:)